OpenStudy (anonymous):

find the lines that are tangent and normal to the curve at the given point: x^2 + xy - y^2 = 1, (2,3)

OpenStudy (anonymous):

looks like an implicit diff question you know how to do that?

OpenStudy (anonymous):

Kind of

OpenStudy (anonymous):

for the middle one you need the product rule that is the only tricky part so remember

OpenStudy (anonymous):

\[x^2+xy-y^2=1\\ 2x+xy'+y-2yy'=0\]

OpenStudy (anonymous):

now you need \(y'\) at \((2,3)\) you have a choice you can solve for \(y'\) using algebra, then plug in \((2,3)\) or you can plug in \(x=2,y=3\) first and solve for \(y'\) that way

OpenStudy (anonymous):

y' should equal -3/2, right?

OpenStudy (anonymous):

i don't know, i didn't do it you want me to check?

OpenStudy (anonymous):

I guess. I just algebra-d the relation out.

OpenStudy (anonymous):

\[4+2y'+3-6y'=0\] would be my first step

OpenStudy (anonymous):

I didn't do that.

OpenStudy (anonymous):

i get \(\frac{7}{4}\) we can try it the other way, really makes no difference

OpenStudy (anonymous):

Nope. I got it wrong. Again

OpenStudy (anonymous):

\[2x+xy'+y-2yy'=0\] \[xy'-2yy'=-2x-y\] \[y'(x-2y)=-2x-y\] \[y'=\frac{-2x-y}{x-2y}\] or \[y'=\frac{2x+y}{2y-x}\]

OpenStudy (anonymous):

then plug in the numbers, i also get \(\frac{7}{4}\)

OpenStudy (anonymous):

Ok, well I know 7/4. Now what do I do?

OpenStudy (anonymous):

use the point slope formula to find the equation of the line though \((2,3)\) with slope \(\frac{7}{4}\)

OpenStudy (anonymous):

that is the tangent line the "normal" line is perpendicular, i.e. the slope is \(-\frac47}{7}\)

OpenStudy (anonymous):

oops \[-\frac{4}{7}\]

OpenStudy (anonymous):

so it's y-3 = 7/4(x-2) y-3 = -4/7(x-2)